Find the inverse function of f(x)=4x+12

1
Expert's answer
2021-11-01T19:51:02-0400

"F(x)=4x+2"

We Isolate X from the expression

Subtract 2 from both sides

"F(x)-2=4x+2-2\\
F(x)-2=4x"

Divide both sides by 4

"X=\frac{F(X)-2}{4}=X"

Replace X by G(X) and F (X) with x

"G(X)=\frac{X-2}{4}"

The inverse of the function is

"G(X)=\frac{X-2}{4}"